# Lattice-induced sound trapping in biperiodic metasurfaces of acoustic resonators

### v3

@@ Abstract
A referential example of a physical system that supports bound states in the continuum (BICs) with an infinite quality factor ($Q$-factor) is a subwavelength lattice of discrete scatterers (resonators) whose response can be significantly modified by exploiting lattice interactions. In this work, we explore the multipole interference mechanism for realizing accidental acoustic BICs (trapped modes) at the $\Gamma$-point (in-plane Bloch wave vector $\mathbf{k}_{\parallel}=\bm{0}$) in biperiodic lattices of acoustic resonators with one resonator per unit cell. To do so, we expand the pressure field from the lattice into a series of scalar zonal ($m=0$) spherical multipoles, carried by a normally incident plane wave, and formulate analytical conditions on the resonator’s multipole moments under which an eigenmode becomes a BIC. The conditions allow us to determine the lattice constant and frequency values that enable the formation of the axisymmetric BIC due to the destructive interference of radiation from zonal multipole moments of a certain parity, although each moment radiates individually. By employing the T-matrix method for acoustic metasurfaces, we numerically investigate the BIC resonance in various structures, including finite arrays, and also the transformation of such resonances into high-$Q$ quasi-BIC regimes, which can be excited by a plane wave at normal incidence.
@@ I Introduction
The physics of discrete electromagnetic or acoustic scatterers arranged in periodic lattices, metasurfaces, attracts a lot of attention because their electromagnetic or acoustic spectra can support additional collective resonances inaccessible with single scatterers [1, 2, 3, 4, 5]. Interactions between an infinite number of scatterers within a lattice can lead to the formation of lattice eigenmodes and significantly modify the lattice response compared to that of isolated single scatterers or finite ensembles [6, 7, 8]. Among all these eigenmodes, trapped modes or bound states in the continuum (BICs) are of particular interest in photonics and acoustics due to their unique characteristics [9, 10, 11, 12, 13, 14]. A genuine BIC has a frequency embedded in the continuum of radiating modes, but an infinite quality factor and no radiative losses [15]. This leads to a confinement (trapping) of the mode energy close to the system or inside it [16].
@@
The formation of BICs has been demonstrated in photonic and acoustic systems using various mechanisms [17, 18, 19, 20, 21, 22, 23]. One of these mechanisms involves coupling between discrete scatterers (particles) arranged in subwavelength lattices (so-called metasurfaces). The electromagnetic (in the case of photonic or microwave systems) or acoustic (in the case of acoustic systems) lattice-induced coupling can lead to the formation of a nonradiant eigenmode if the lattice constant is smaller than the wavelength in the background. In such a situation, a lattice eigenmode can radiate only in two directions determined by its in-plane Bloch wave vector $\mathbf{k}_{\parallel}$. If the far field of the mode vanishes in these two directions, the mode is commonly referred to as a symmetry-protected (or $\Gamma$-point) BIC for $\mathbf{k}_{\parallel}=\bm{0}$, and as an accidental (or parametric) BIC for $\mathbf{k}_{\parallel}\neq\bm{0}$ [24]. However, this distinction is not exclusive, and accidental BICs have also been shown to be accessible in photonic systems when $\mathbf{k}_{\parallel}=\bm{0}$ [25, 26, 27, 28]. One of the approaches involves the formation of a nonradiant state due to the desctructive interference of multipoles with the same parity (symmetry with respect to the inversion $\mathbf{r}\to-\mathbf{r}$) [28].
@@
A comprehensive description of BICs in subwavelength lattices can be obtained using the multipole decomposition method [24], which has been successfully applied many times to describe lattice resonant effects in photonics and acoustics [28, 29, 30, 31, 32, 33]. This method is based on the expansion of an arbitrary field into mutually orthogonal functions with unique radiation patterns called multipoles, which provide a set of basic solutions to the Helmholtz equation [34]. Here, under multipoles, we assume spherical waves. It is important to emphasize that the multipole decomposition differs between electromagnetics and acoustics, as an electromagnetic wave is vectorial and transverse. In contrast, an acoustic pressure wave is a scalar and longitudinal wave. In the first case, the decomposition involves the vector electric and magnetic multipoles (spherical waves) of degree $\ell\geq 1$. In contrast, in the second case, it only considers the scalar multipoles with $\ell\geq 0$. The fact that photonic systems possess simultaneously two types of multipoles (electric and magnetic) and resonances provides many degrees of freedom to tailor their optical response based on interactions [35, 36, 37, 38]. Although acoustic scatterers only possess one type of multipoles (scalar), the presence of the monopole ($\ell=0$), which can interfere with the dipole ($\ell=1$) and higher-degree multipoles, gives rise to acoustics-specific effects, including, e.g., the acoustic Kerker effect [39, 40] and the lateral recoil force [41].
@@
The distinction outlined above necessitates different approaches to realize a BIC response in photonic and acoustic systems. The multipolar content of a symmetry-protected BIC (s-BIC) can contain only multipoles with $m\neq\pm 1$ that do not radiate in the normal direction with respect to the lattice [24]. The simplest but feasible example of an electromagnetic s-BIC is a lattice of vertically oriented electric or magnetic dipoles ($\ell=1,m=0$) [24, 12]. Indeed, such a dipole does not radiate transverse waves in the normal (vertical) direction. This makes the entire lattice nonradiant if the dipole moments of the particles have the same amplitude and phase, i.e., $\mathbf{k}_{\parallel}=\bm{0}$. Moreover, such a mode can be excited under normal incidence only by breaking the particle symmetry [42, 43, 44, 45, 46], so that it becomes $C_{1v}$ [47]. This can be achieved, for example, if the sphere is replaced by a nonequilateral prism. The symmetry breaking enables bianisotropy-induced coupling in isolated photonic scatterers between the nonradiant electric dipole from the s-BIC with $m=0$ and the magnetic dipole of the external wave with $m=\pm 1$, or vice versa [48, 49]. In this case, the state is referred to as a quasi-s-BIC to emphasize a finite but large $Q$-factor. A quasi-BIC can be of even more practical importance than a genuine BIC, since it can be excited by an external plane wave at normal incidence [50, 51].
@@
In acoustics, a normally incident pressure plane wave is a superposition of scalar multipoles that always radiate in the normal direction – the so-called zonal multipoles, which have an order of $m=0$. Although scalar multipoles with $m\neq 0$ do not radiate in the normal direction and therefore can form acoustic s-BICs, it can be hindered, if not impossible, to efficiently excite an acoustic quasi-s-BIC with a high $Q$-factor under normal incidence if we do not break the particle rotational symmetry [52]. Recently, it was shown in Ref. 28 that lattice-induced coupling between multipoles of the same parity (and $m=\pm 1$) can enable optical accidental BICs (a-BICs) at the $\Gamma$-point of a metasurface of spherical resonators, while each individual multipole can be excited by an external wave, resulting in quasi-a-BICs with high-$Q$ resonances under normal incidence without breaking the metasurface symmetry. Therefore, in this work, we extend and explore this approach for acoustic lattices irradiated by a normally incident pressure wave [see Fig. 1(a)], showing its consistency when mapped onto a different wave propagation platform. We demonstrate that subwavelength 2D lattices of particles (acoustic resonators) can also support accidental acoustic BICs and quasi-BICs for $\mathbf{k}_{\parallel}=\bm{0}$, emerging due to the monopole-quadrupole or dipole-octupole coupling that leads to suppression of acoustic radiation in the normal direction [see Fig. 1(b)]. For this, in Sec. II, we present a comprehensive theoretical model, based on the T-matrix method and multipole decomposition, to describe the acoustic response of 2D lattices as well as their eigenmodes. This model enables us to obtain in Sec. III analytical conditions under which the acoustic axisymmetric BICs emerge. Moreover, in Sec. III, we model them in infinite arrays of spherical resonators and discuss their transformation into quasi-BICs as the lattice constant and frequency are tuned. Finally, in Sec. IV, we investigate the quasi-BIC regime under realistic conditions such as the presence of material losses in the particles or a substrate, as well as finite sizes of the array.
@@
Figure 1: (a) Sketch of a biperiodic lattice of spherical particles (resonators) with a subwavelength lattice constant of $L$; the structure supports acoustic bound states in the continuum (trapped modes) at the $\Gamma$-point (normal incidence). (b) Formation of the acoustic BIC thank to the destructive interference of scalar zonal multipoles with the same parity (with respect to $\mathbf{r}\to-\mathbf{r}$). The diagrams illustrate the intensity of the pressure field in the far-field region (i.e., $r\gg\lambda_{\rm b}$) generated by a unit cell. The color represents the pressure field phase, while the radius indicates its amplitude.
@@ II.1 Basic equations
In linear acoustics, the propagation of a monochromatic acoustic wave in a medium with compressibility $\beta$ and mass density $\rho$ is described by pressure $p(\omega;\mathbf{r})$ and velocity $\mathbf{v}(\omega;\mathbf{r})$ that obey the laws of mass and linear momentum conservation [53, 54]
@@
By substituting the second equation into the first one, we see that the pressure field obeys the acoustic wave equation
@@
where $c=1/\sqrt{\beta\rho}$ is the speed of sound. Formally, (2) presents the homogeneous Helmholtz equation with wavenumber $k=\omega/c$.
II.2 T-matrix approach to multiple scattering problems
@@ II.2.1 Single resonator
An efficient approach to calculating the scattered field in spherical coordinates, which involves expanding $p_{\rm inc}(\omega;\mathbf{r})$ and $p_{\rm sca}(\omega;\mathbf{r})$ into a series of fundamental solutions of Eq. (2), was laid out by Lord Rayleigh in his seminal work, The Theory of Sound (see Ch. 17, [55]). In today’s notation, such fundamental solutions, the scalar spherical waves or simply multipoles $\Psi^{(n)}_{\ell,m}(\omega;\mathbf{r})$ with $\ell=0,1,2,$... and $m=-\ell,-\ell+1,...,\ell-1,\ell$ being the degree and order [56], are commonly used to model acoustic sources and scatterers. The definition of $\Psi^{(n)}_{\ell,m}(\omega;\mathbf{r})$ is given in Appendix A. In the case of the acoustic scattering by an individual particle (resonator), the series takes the following form (we assume that the series converges to the actual fields outside the smallest sphere encapsulating the entire scatterer)
@@
where $\sum_{\ell,m}\equiv\sum_{\ell=0}^{+\infty}\sum_{m=-\ell}^{\ell}$. The arguments of the expansion coefficients include the origin of the multipole expansion, denoted here as $\mathbf{r}_{0}$. Following Ref. 57, we assume that $\mathbf{r}_{0}$ coincides with the center of mass of the scatterer and the origin of the coordinate system to omit the dependence on $\mathbf{r}_{0}$ further. To highlight the dependence of the coefficients on the frequency $\omega$ in the general case, we explicitly write it as an argument. However, for a pressure plane wave, the coefficients $b_{\ell,m}(\omega)$ do not depend on the frequency $\omega$, but depend on the direction of incidence (see Appendix B).
@@
where $\mathbf{a}(\omega)$ and $\mathbf{b}(\omega)$ are the column vectors of the corresponding coefficients arranged in ascending order of $\ell=0,1,2,$... and $m=-\ell,-\ell+1,$$\ell$; and $\mathbf{T}(\omega)$ is the T-matrix after a truncation of the series at $\ell_{\rm max}$ (large enough to reach convergence). The T-matrix contains the necessary and sufficient “information” about the scatterer to describe the scattering of an arbitrary field $p_{\rm inc}(\omega;\mathbf{r})$ at a given frequency [58, 59]. The elements of the T-matrix are calculated analytically for the spherical scatterers (see Ref. 32 and also Appendix C) and numerically for arbitrarily shaped scatterers [60].
II.2.2 Periodic lattice of resonators
The T-matrix method also provides an analytical, clear, and efficient treatment of the scattering of a plane wave by periodic lattices [61, 31]. We assume a lattice in the $xy$ plane with a simple unit cell, i.e. one acoustic scatterer per unit cell [see Fig. 1(a)]. First, we notice that, because of the lattice translation symmetry, the multipole moments of the unit cell at $\mathbf{R}$ are linked to those of the reference unit cell at $\bm{0}$ via Bloch’s theorem,
@@ II.2.2 Periodic lattice of resonators
where the in-plane wave vector $\mathbf{k}_{\parallel}$ determines the phase shift between the unit cells. In the presence of an incident plane wave, $\mathbf{k}_{\parallel}$ is determined by the in-plane component of its wave vector, i.e. $\mathbf{k}=\mathbf{k}_{\parallel}\pm\sqrt{k^{2}-k_{\parallel}^{2}}\>\widehat{\mathbf{z}}$, where $\widehat{\mathbf{z}}$ is the unit out-of-plane vector directed along the $z$ axis of the Cartesian coordinate system in Fig. 1(a). In the case of an eigenmode, $\mathbf{k}_{\parallel}$ is its Bloch wave vector from the first Brillouin zone. A new subscript “eff” is introduced to distinguish the multipole moments of the scatterer embedded in the lattice (11) from the moments ${\bf a}(\omega)$ of the isolated scatterer from (7). This difference arises because, when a scatterer is placed on a lattice, the scatterers interact with each other. Hence, the unit cell with one scatterer placed at $\mathbf{r}_{0}=\bm{0}$ (reference unit cell) experiences (i) the incident field $p_{\rm inc}(\omega;\mathbf{r})$, and (ii) the secondary fields, rescattered by all the other scatterers located at the lattice nodes $\mathbf{R}\neq\bm{0}$ and determined by the multipole coefficients $\mathbf{a}^{\rm eff}_{\mathbf{R}}$. The lattice-mediated interaction between the unit cells can be accommodated within the T-matrix method by multiplication with the lattice sum matrix $\bm{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})$ in Eq. (7), where denotes that the reference unit cell is excluded from the sum [see Eq. (42)]. Hence, the multipole coefficients of the reference unit cell can be computed by solving the following equation (see Appendix D for details)
$\mathbf{a}^{\rm eff}=\mathbf{T}(\omega)\left[\mathbf{b}+\bm{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})\mathbf{a}^{\rm eff}\right]\,$, (9)
@@
Equation (10) can be used to compute the acoustic response of the lattice to an external wave (e.g., plane wave). In this case, a formal solution to Eq. (10) is given by
@@
where we have used (7). To recall the dependence of the vector of effective multipole coefficients on $\omega$ and $\mathbf{k}_{\parallel}$, we write $\mathbf{a}^{\rm eff}$ as $\mathbf{a}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$ if such a highlight is needed.
@@
Note that while vector $\mathbf{a}$ of the single scatterer in (11) depends on $\omega$, it also depends on $\mathbf{k}_{\parallel}$, but only because the expansion of a plane wave into spherical multipoles $\mathbf{b}$ depends on $\mathbf{k}$ (see Appendix B). In contrast to the single-particle T-matrix $\mathbf{T}(\omega)$, which is calculated once at a given frequency and used for an arbitrary incident wave, the matrix $\mathbf{T}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ in (11) must be recomputed for all the desired values of $\mathbf{k}_{\parallel}$ at a given frequency.
@@
It also follows from (11) that the coupling between an infinite number of scatterers within the lattice leads to the renormalization of their acoustic multipole response $\mathbf{a}^{\rm eff}=\left[\mathbf{I}-\mathbf{T}(\omega)\mathbf{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})\right]^{-1}\mathbf{a}$. The scattered (radiated) pressure field in a lattice is a superposition of the fields emitted by an infinite number of particles on the lattice [see, e.g., Eq. (13)]. When the determinant of $\mathbf{T}_{\rm eff}^{-1}$ in (10) approaches zero, we observe resonant features in the lattice spectrum. A detailed discussion of these features, related to the lattice eigenmodes [61, 33, 62], is given in Section III.
II.3 Selection rules for multipole coupling in 2D lattices

### v4

@@ Abstract
A referential example of a physical system that supports bound states in the continuum (BICs) with an infinite quality factor ($Q$ factor) is a metasurface of discrete scatterers (resonators), whose response can be significantly modified by exploiting lattice interactions. In this work, we explore the multipole-interference mechanism for realizing accidental acoustic BICs (trapped modes) at $\Gamma$-point (in-plane Bloch wave vector $\mathbf{k}_{\parallel}=\mathbf{0}$) of biperiodic metasurfaces of acoustic resonators with one resonator per unit cell. To do so, we expand the pressure field from the metasurface into a series of scalar zonal ($m=0$) spherical multipoles, carried by a normally incident plane wave, and formulate analytical conditions on the resonator multipole moments under which an eigenmode becomes a BIC. The conditions enable us to determine the lattice constant and frequency values that facilitate the formation of an axisymmetric BIC with a specific parity, resulting from destructive interference between zonal multipoles of the same parity, despite each moment radiating individually. By employing the T-matrix method for acoustic metasurfaces, we numerically investigate the BIC resonance in various structures, including finite arrays, and also the transformation of such resonances into high-$Q$ quasi-BIC regimes, which can be excited by a plane wave at normal incidence.
@@ I Introduction
The physics of discrete electromagnetic or acoustic scatterers arranged in periodic lattices attracts significant attention because their electromagnetic or acoustic spectra can support additional collective resonances inaccessible with single scatterers [1, 2, 3, 4, 5]. Interactions between an infinite number of scatterers within a lattice can lead to the formation of lattice eigenmodes and significantly modify the lattice response compared to that of isolated scatterers or finite ensembles [6, 7, 8]. Among these eigenmodes, trapped modes or bound states in the continuum (BICs) are of particular interest in photonics and acoustics due to their unique characteristics [9, 10, 11, 12, 13, 14]. A genuine BIC has a frequency embedded in the continuum of radiating modes, but an infinite quality factor and no radiative losses [15]. This leads to confinement (trapping) of the mode energy near or inside the system [16].
@@
The formation of BICs has been demonstrated in photonic and acoustic systems using various mechanisms [17, 18, 19, 20, 21, 22, 23, 24]. One of these mechanisms involves coupling between discrete scatterers (particles) arranged in subwavelength biperiodic lattices, also referred to as metasurfaces. The electromagnetic (in photonic or microwave systems) or acoustic (in acoustic systems) lattice-induced coupling can lead to the formation of a nonradiant eigenmode if the lattice constant is smaller than the wavelength in the background. In such a situation, a lattice eigenmode can radiate only in directions determined by its in-plane Bloch wave vector $\mathbf{k}_{\parallel}$. If the far field of the mode vanishes in these directions, the mode is commonly referred to as a symmetry-protected (or $\Gamma$-point) BIC for $\mathbf{k}_{\parallel}=\bm{0}$, and as an accidental (or parametric) BIC for $\mathbf{k}_{\parallel}\neq\bm{0}$ [25]. However, this distinction is not exclusive, and accidental BICs have also been shown to occur in photonic systems at $\mathbf{k}_{\parallel}=\bm{0}$ [26, 27, 28, 29]. One approach involves forming a nonradiant state through destructive interference of multipoles with the same parity (symmetry with respect to inversion $\mathbf{r}\to-\mathbf{r}$) [29].
@@
A comprehensive description of BICs in metasurfaces can be obtained using the multipole decomposition method [25], which has been successfully applied to describe lattice resonant effects in photonics and acoustics [29, 30, 31, 32, 33, 34]. This method expands an arbitrary field into mutually orthogonal functions with unique radiation patterns, called multipoles, providing a set of basic solutions to the Helmholtz equation [35]. Here, multipoles are assumed to be spherical waves. It is important to note that the multipole decomposition differs between electromagnetics and acoustics: electromagnetic waves are vectorial and transverse, whereas acoustic pressure waves are scalar and longitudinal. In electromagnetics, the decomposition involves vector electric and magnetic multipoles of degree $\ell\geq 1$, while in acoustics, it only involves scalar multipoles of degree $\ell\geq 0$. Thus, photonic systems possess two types of multipoles and resonances, providing multiple degrees of freedom to tailor the optical response [36, 37, 38, 39]. Although acoustic scatterers possess only scalar multipoles, the presence of the monopole ($\ell=0$), which can interfere with the dipole ($\ell=1$) and higher-degree multipoles, gives rise to acoustics-specific effects, such as the acoustic Kerker effect [40, 41] and the lateral recoil force [42].
@@
The distinction outlined above necessitates different approaches to realize a BIC response in photonic and acoustic systems. The multipolar content of a symmetry-protected BIC (s-BIC) can contain only multipoles with $m\neq\pm 1$, which do not radiate in the normal direction with respect to the lattice [25]. The simplest example of an electromagnetic s-BIC is a lattice of vertically oriented electric or magnetic dipoles ($\ell=1,m=0$) [25, 12]. Such a dipole does not radiate transverse waves in the normal direction, making the entire lattice nonradiant if the dipole moments of all particles have the same amplitude and phase, i.e., $\mathbf{k}_{\parallel}=\bm{0}$. Moreover, such a mode can be excited under normal incidence only by breaking the particle symmetry [43, 44, 45, 46, 47], which breaks inversion and rotational symmetry with respect to the $z$ axis [48]. This can be achieved, for example, by replacing a sphere with a non-equilateral prism. This symmetry breaking enables bianisotropy-induced coupling between the nonradiant electric dipole from the s-BIC with $m=0$ and the magnetic dipole of the external wave with $m=\pm 1$, or vice versa [49, 50]. This state is referred to as a quasi-s-BIC to emphasize its finite but large $Q$ factor. A quasi-BIC can be more practically important than a genuine BIC, as it can be excited by an external plane wave at normal incidence [51, 52].
@@
In acoustics, a normally incident pressure plane wave is a superposition of scalar multipoles that always radiate in the normal direction – so-called zonal multipoles with $m=0$. Although scalar multipoles with $m\neq 0$ do not radiate in the normal direction and can form acoustic s-BICs, it can be hindered, if not impossible, to efficiently excite an acoustic quasi-s-BIC with a high $Q$-factor under normal incidence without breaking the particle rotational symmetry [53]. Recently, Allayarov et al. showed that lattice-induced coupling between multipoles of the same parity (and $m=\pm 1$) can enable accidental optical BICs (a-BICs) at the $\Gamma$-point of a metasurface of spherical resonators [29]. In this case, each multipole can be excited by an external wave, resulting in quasi-a-BICs with high-$Q$ resonances under normal incidence without breaking metasurface symmetry.
Therefore, in this work, we extend and explore this approach for acoustic metasurfaces excited by a normally incident pressure wave [see Fig. 1(a)], showing its consistency when mapped onto a different wave propagation platform. We demonstrate that metasurfaces with inversion symmetry can also support accidental acoustic BICs and quasi-BICs for $\mathbf{k}_{\parallel}=\bm{0}$, arising from monopole–quadrupole or dipole–octupole coupling that suppresses acoustic radiation in the normal direction [see Fig. 1(b)]. In Sec. II, we present a comprehensive theoretical model to describe the acoustic response of metasurfaces and their eigenmodes. This model enables us to derive in Sec. III, analytical conditions under which axisymmetric acoustic BICs emerge. Moreover, in Sec. III, we model them in infinite arrays of spherical resonators and discuss their transformation into quasi-BICs as the lattice constant and frequency are tuned. Finally, in Sec. IV, we investigate the quasi-BIC regime under realistic conditions, such as the presence of material losses in the particles or a substrate, as well as finite array sizes.
@@
Figure 1: (a) Sketch of a biperiodic lattice of spherical particles (resonators) with a subwavelength lattice constant of $L$; the structure supports accidental acoustic bound states in the continuum (trapped modes) with a certain parity at the $\Gamma$ point (normal incidence). (b) Formation of the acoustic BICs owing to destructive interference between scalar zonal multipoles with the same parity (with respect to $\mathbf{r}\to-\mathbf{r}$). The diagrams illustrate the intensity of the pressure field in the far-field region, $r\gg\lambda_{\rm b}$, generated by a unit cell. The color represents the pressure field phase, while the radius indicates its amplitude.
@@ II.1 Basic equations
In linear acoustics, the propagation of a monochromatic acoustic wave in a medium with compressibility $\beta$ and mass density $\rho$ is described by pressure $p(\omega;\mathbf{r})$ and velocity $\mathbf{v}(\omega;\mathbf{r})$ that obey the laws of mass and linear momentum conservation [54, 55]:
@@
By substituting the second equation into the first one, we see that the pressure field obeys the acoustic wave equation,
@@
where $c=1/\sqrt{\beta\rho}$ is the speed of sound. Formally, Eq. (2) presents the homogeneous Helmholtz equation with wavenumber $k=\omega/c$.
II.2 T-matrix approach to scattering of acoustic waves
@@ II.2.1 Single resonator
An efficient approach to calculate the scattered field in spherical coordinates, which involves expanding $p_{\rm inc}(\omega;\mathbf{r})$ and $p_{\rm sca}(\omega;\mathbf{r})$ into a series of fundamental solutions of Eq. (2), was laid out by Lord Rayleigh in his seminal work, The Theory of Sound (see Chap. 17 of Ref. 56). In today‘s notation, such fundamental solutions, the scalar spherical waves or simply multipoles $\Psi^{(n)}_{\ell,m}(\omega;\mathbf{r})$ with $\ell=0,1,2,$... and $m=-\ell,-\ell+1,...,\ell-1,\ell$ being the degree and order [57], are commonly used to model acoustic sources and scatterers. The definition of $\Psi^{(n)}_{\ell,m}(\omega;\mathbf{r})$ is given in Appendix A. In the case of the acoustic scattering by an individual particle (resonator), the series takes the following form (we assume that the series converges to the actual fields outside the smallest sphere encapsulating the entire scatterer):
@@
where $\sum_{\ell,m}\equiv\sum_{\ell=0}^{+\infty}\sum_{m=-\ell}^{\ell}$. The arguments of the expansion coefficients include the origin of the multipole expansion, denoted here as $\mathbf{r}_{0}$. Following Ref. 58, we assume that $\mathbf{r}_{0}$ coincides with the center of mass of the scatterer and the origin of the coordinate system to omit the dependence on $\mathbf{r}_{0}$ further. To highlight the dependence of the coefficients on the frequency $\omega$ in the general case, we explicitly write it as an argument. However, for a pressure plane wave, the coefficients $b_{\ell,m}(\omega)$ do not depend on the frequency $\omega$, but depend on the direction of incidence (see Appendix B).
@@
where $\mathbf{a}(\omega)$ and $\mathbf{b}(\omega)$ are the column vectors of the corresponding coefficients arranged in ascending order of $\ell=0,1,2,$... and $m=-\ell,-\ell+1,$$\ell$, and $\mathbf{T}(\omega)$ is the T-matrix after a truncation of the series at $\ell_{\rm max}$ (large enough to reach convergence). The T-matrix contains the necessary and sufficient “information” about the scatterer to describe the scattering of an arbitrary field $p_{\rm inc}(\omega;\mathbf{r})$ at a given frequency [59, 60]. The elements of the T-matrix are calculated analytically for the spherical scatterers (see Ref. 33 and also Appendix C) and numerically for arbitrarily shaped scatterers [61].
II.2.2 Biperiodic metasurface of resonators
The T-matrix method also provides an analytical, clear, and efficient treatment of the scattering of a plane wave by periodic metasurfaces [62, 32]. We assume a lattice in the $xy$ plane with a simple unit cell, i.e., one acoustic scatterer per unit cell [see Fig. 1(a)]. First, we notice that, because of the translation symmetry, the multipole moments of the unit cell at $\mathbf{R}$ are linked to those of the reference unit cell at $\bm{0}$ via Bloch’s theorem,
@@ II.2.2 Biperiodic metasurface of resonators
where the in-plane wave vector $\mathbf{k}_{\parallel}$ determines the phase shift between the unit cells. In the presence of an incident plane wave, $\mathbf{k}_{\parallel}$ is determined by the in-plane component of its wave vector, i.e. $\mathbf{k}=\mathbf{k}_{\parallel}\pm\sqrt{k^{2}-k_{\parallel}^{2}}\>\widehat{\mathbf{z}}$, where $\widehat{\mathbf{z}}$ is the unit out-of-plane vector directed along the $z$ axis of the Cartesian coordinate system in Fig. 1(a). In the case of an eigenmode, $\mathbf{k}_{\parallel}$ is its Bloch wave vector from the first Brillouin zone. A subscript “eff” is introduced to distinguish the multipole moments of the scatterer embedded in the lattice (11) from the moments ${\bf a}(\omega)$ of the isolated scatterer from Eq. (7). This difference arises because, when a scatterer is placed on a lattice, the scatterers interact with each other. Hence, the unit cell with one scatterer placed at $\mathbf{r}_{0}=\bm{0}$ (reference unit cell) experiences (i) the incident field $p_{\rm inc}(\omega;\mathbf{r})$, and (ii) the secondary fields, rescattered by all the other scatterers located at the lattice nodes $\mathbf{R}\neq\bm{0}$ and determined by the multipole coefficients $\mathbf{a}^{\rm eff}_{\mathbf{R}}$. The lattice-induced interaction between the unit cells can be accommodated within the T-matrix method by multiplication with the lattice sum matrix $\bm{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})$ in Eq. (7), where denotes that the reference unit cell is excluded from the sum [see Eq. (43)]. Hence, the multipole coefficients of the reference unit cell can be computed by solving the following equation (see Appendix D for details)
$\mathbf{a}^{\rm eff}=\mathbf{T}(\omega)\left[\mathbf{b}+\bm{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})\mathbf{a}^{\rm eff}\right]$ (9)
@@
Equation (10) can be used to compute the acoustic response of a metasurface to an external wave (e.g., plane wave). In this case, a formal solution to Eq. (10) is given by
@@
where we have used Eq. (7). To recall the dependence of the vector of effective multipole coefficients on $\omega$ and $\mathbf{k}_{\parallel}$, we write $\mathbf{a}^{\rm eff}$ as $\mathbf{a}^{\rm eff}(\omega,\mathbf{k}_{\parallel})$ if such a highlight is needed.
@@
Note that while vector $\mathbf{a}$ of the single scatterer in Eq. (11) depends on $\omega$, it also depends on $\mathbf{k}_{\parallel}$, but only because the expansion of a plane wave into spherical multipoles $\mathbf{b}$ depends on $\mathbf{k}$ (see Appendix B). In contrast to the single-particle T-matrix $\mathbf{T}(\omega)$, which is calculated once at a given frequency and used for an arbitrary incident wave, the matrix $\mathbf{T}_{\rm eff}(\omega,\mathbf{k}_{\parallel})$ in Eq. (11) must be recomputed for all the desired values of $\mathbf{k}_{\parallel}$ at a given frequency.
@@
It also follows from Eq. (11) that the coupling between an infinite number of scatterers within a metasurface leads to the renormalization of their acoustic multipole response $\mathbf{a}^{\rm eff}=\left[\mathbf{I}-\mathbf{T}(\omega)\mathbf{\Sigma}^{\prime}(\omega,\mathbf{k}_{\parallel})\right]^{-1}\mathbf{a}$. The scattered (radiated) pressure field for a metasurface is a superposition of the fields emitted by an infinite number of particles on the lattice [see, e.g., Eq. (13)]. When the determinant of $\mathbf{T}_{\rm eff}^{-1}$ in Eq. (10) approaches zero, we observe resonant features in the spectrum of the metasurface. A detailed discussion of these features, related to the lattice eigenmodes [62, 34, 63], is given in Sec. III.
II.3 Selection rules for multipole coupling in metasurfaces